An airplane flies 440 miles with the wind and 340 against the wind in the same length of time. If the speed of the wind is 20, what is the speed of the airplane in still air?
Answer by mananth(16949)

plane speed x mph
wind speed 20 mph
..
against wind x - 20 mph
with wind x + 20 mph

Distance against wind 340
Distance with wind 440
Time = same
340 /(x- 20 )= 440 /(x+ 20 )
340 *(x+ 20 )= 440 *(x- 20 )
340 x + 6800 = 440 x -8800
440 x -340 x = 8800 + 6800
100 x = 15600
/ 100
x= 156 mph
speed in still air = 156 mph